• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Mitra s

    Log entry time 16:01:46 on March27,2014

    Entry number 387812

    keyword=Day Shift Summary

    Thursday 3/27/2014
    Day Shift Summary:
    --------------------------------------------
    SL: Mitra H. Shabestari
    TO: Nicholas Hirlinger Saylor
    3rd: Camille Desnault
    --------------------------------------------

    8:00---------- No beam.
    8:15---------- Called MCC and we should get beam in about 1 hour.
    8:30---------- Deadtime reading for left arm is wrong. It need to be corrected when we access the hall.
    9:15---------- RC decided to change the run plan.
    Put RHRS sieve in and changed the momentum from 1GeV to 2GeV.
    LHRS momentum is at 3GeV.
    9:50---------- Josh called to let us know that target group were going to try to fix the fan problem remotely, so we should ignore target alarms.
    10:30---------- No beam yet. Contacted MCC and we should get beam in one hour.
    10:45---------- Target group are done. Dave called; it is safe to put LH2 target in
    only with 2X2 raster and maximum current of 5uA.
    11:00---------- RHRS at 2GeV and NMR is locked.
    15:16---------- No Beam. According to MCC there won't be beam for a couple of hours.
    16:00---------- Day shift ended with no Beam.


    --------------------* NOTE *--------------------

    Issues need to be taken care of:

    -- The reading for the Loop1 on the target panel on top of the target console is wrong. While target is at 19 Kelvins it reads ~100 Kelvins.

    -- Deadtime reading on the left arm is meaningless. Has to be fixed when we get access to the hall.